2xy′′ + (x+1)y ′ +3y = 0; Find the series solution of the following equations near x = 0:


1
Expert's answer
2021-02-02T05:45:21-0500

We will substitute solution of the form "y=\sum_{k=0}^{+\infty}A_kx^k" in the equation. We receive: "y'=\sum_{k=1}^{+\infty}kA_kx^{k-1}" and "y''=\sum_{k=2}^{+\infty}k(k-1)A_kx^{k-2}". After subsitution we receive: "2x\sum_{k=2}^{+\infty}k(k-1)A_kx^{k-2}+(x+1)\sum_{k=1}^{+\infty}kA_kx^{k-1}+3\sum_{k=0}^{+\infty}A_kx^k=0".

We simplify and receive:

"2\sum_{k=2}^{+\infty}k(k-1)A_kx^{k-1}+\sum_{k=1}^{+\infty}kA_kx^{k}+\sum_{k=1}^{+\infty}kA_kx^{k-1}+3\sum_{k=0}^{+\infty}A_kx^k=0".

We change summation indices and get:

"2\sum_{j=1}^{+\infty}(j+1)jA_{j+1}x^{j}+\sum_{k=1}^{+\infty}kA_kx^{k}+\sum_{j=0}^{+\infty}(j+1)A_{j+1}x^{j}+3\sum_{k=0}^{+\infty}A_kx^k=0".

Finally, we get: "A_1+3A_0+2\sum_{k=1}^{+\infty}(k+1)kA_{k+1}x^{k}+\sum_{k=1}^{+\infty}kA_kx^{k}+\sum_{k=1}^{+\infty}(k+1)A_{k+1}x^{k}+3\sum_{k=1}^{+\infty}A_kx^k=0"

We obtain:

"A_1+3A_0+\sum_{k=1}^{+\infty}((k+1)(2k+1)A_{k+1}+(k+3)A_k)x^k=0"

Thus, we can choose an arbitrary "A_0" and then set "A_1=-3A_0", "6A_2+4A_1=0", "15A_3+5A_2=0"

As a result, we received the recurrence formula for the coefficients given by: "A_1=-3A_0," "A_2=-\frac23A_1", "A_3=-\frac13A_2,"... In general, for "k\geq1" we get: "(k+1)(2k+1)A_{k+1}+(k+3)A_k=0".